OpenStudy (anonymous):

what is the sum of these mixed numbers ? 3 1/6 + 3 1/2 = what i know that the outside number is 6

OpenStudy (anonymous):

3 1/6 = 19/6 3 1/2 = 7/2 \[\left(\begin{matrix}7 \\ 2\end{matrix}\right)\rightarrow \left(\begin{matrix}21 \\ 6\end{matrix}\right) + \left(\begin{matrix}19 \\ 6\end{matrix}\right)= \left(\begin{matrix}40 \\ 6\end{matrix}\right)\] How many times foes 6 go into 40? 6 times. What's the remainder? 4. Answer: 6 4/6 or 6 2/3

OpenStudy (anonymous):

get it?